OpenStudy (anonymous):

Let θ = 135°. Find the exact values of tan θ , cos θ & csc θ

OpenStudy (amistre64):

they want youto use the angle addition formulas

OpenStudy (amistre64):

but its really just 45 degrees

OpenStudy (amistre64):

-cos; and sin

OpenStudy (anonymous):

tan 135=-1 cos 135=-0.707 csc 135=1.414

OpenStudy (amistre64):

sin(45) = 1/sqrt(2); csc(45) = sqrt(2) -cos(45) = -sqrt(2)/2 -tan(45) = -1

OpenStudy (amistre64):

decimals are 'exact' values; gotta keep the sqrts

myininaya (myininaya):

sin(135)=sin(180-45)=cos(180)sin(45)-cos(45)sin(180) cos(135)=sin(180-45)=cos(180)cos(45)+sin(180)sin(45) use this

OpenStudy (amistre64):

decimals ARENT exact.... my keyboard hates me

OpenStudy (anonymous):

wait so whats the answer ive got'n like 30 diffedrent replies lol

OpenStudy (anonymous):

vote if my answers r right

OpenStudy (anonymous):

use myininaya's way to do it (your using the complimentary angle_)